how many ounces of a 100% juice concentrate must be added to a 20% juice concentrate to yield 128 ounces of a 50% juice
Answer by ikleyn(52781)

.
how many ounces of a 100% juice concentrate must be added to a 20% juice concentrate to yield 128 ounces of a 50% juice
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Solution 1. Two equations

x +    y = 128,       (1)   (x is the volume of the 100% juice)
x + 0.2y = 0.5*128.   (2)

Simplify

x +    y = 128,       (1)
x + 0.2y =  64.       (2)

To solve it, distract (2) from (1). You will get

0.8y = 128-64  --->  0.8y = 64.

From this point, please complete the solution on your own.


Solution 2. One equation

x + 0.5(128-x) = 0.5*128.    (x is the volume of the 100% juice)

Simplify and solve for x.

There is entire bunch of the lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

It contains many other solved word problems, as well as many other interesting and useful things.